# No Language is an Island: Unifying Chinese and English in Financial Large Language Models, Instruction Data, and Benchmarks

### v1

@@ E Examples of Instruction Prompt in Chinese Task
Determine the sentiment of the sentences in the provided financial news article, identifying whether the sentiment is positive, negative, or neutral. Simply answer ’positive’, ’neutral’, or ’negative’.
Sentiment Matching (FinSM)
判断给出的两个金融文本表达的意思是否相似,你只需要回答是或否。现在判断下面两个句子是否相似:
@@
Determine whether the two provided financial texts convey similar meanings; you only need to answer ’yes’ or ’no’. Now determine if the following two sentences are similar.
News Classification (FinNC)
根据金融报道的内容,判断其属于’中国’、’国际’还是’外国’,只需给出其中一个类别。现在分析下面的报道,回答它所属的类别:
@@
According to the content of the financial report, determine whether it belongs to ’China’, ’International’, or ’Foreign’, only one category needs to be given. Now analyze the following report and answer the category it belongs to:
Negative Judgment (FinNJ)
根据所给的金融新闻及实体,你需要判断所给实体是否含有负面的消息,你只需简单回答’有’或者’无’
@@
Based on the given financial news and entities, you need to determine whether the given entities contain negative messages. You only need to simply answer ’yes’ or ’no’.
Answer Selection (FinAS)
根据所提出的金融问题,从以下四个选项中选择最合适的一个。你的输出应该是:’A’、’B’、’C’ 或’D’。
@@
Based on the financial issue presented, choose the most appropriate one from the following four options. Your output should be: ’A’, ’B’, ’C’, or ’D’.
Stock Prediction (FinSP)
请仔细分析数据和推文,预测2017-10-13时\$trv的收盘价格是上涨还是下跌。请确认是上涨还是下跌, 只回答为’上涨’或者’下跌’。
@@
Please carefully analyze the data and tweets, and predict whether the closing price of \$TRV on 2017-10-13 will go up or down. Please confirm whether it is going up or down. You can answer ’Rise’ or ’Fall’.
Headline Classification (FinHC)
请考虑标题是否讨论了与黄金相关的过去事件。新闻标题是否暗示黄金的过去新闻?您的回答应该为’是’或’否’。
@@
Please consider whether the headline discusses past events related to gold. Does the news headline imply past news about gold? Your answer should be ’Yes’ or ’No’.
Question Answering (FinQA)
你需要分析金融文本,根据内容回答相关问题。如果你对答案感到迷茫,可以回答’无相应参数’。
@@
You need to analyze financial texts and answer related questions based on the content. If you feel confused about the answer, you can reply with ’No relevant parameters’.
Event Detection (FinED)
阅读以下金融领域的公告,判断所有的事件类型及其对应主体,省略数字。请以:’事件类型, 事件主体’的格式回复 。其中事件类型应该在这里面:[’信批违规’,…,’涉嫌欺诈’]
@@
The task is to read a financial announcement, determine all event types and their corresponding entities, and reply in the format ’event type, event subject’. The event types should be among those listed in the brackets [‘credit approval violation’,…,’suspected fraud’].
Entity Recognition (FinER)
在分析中国证券交易委员会备案文件中的句子时,识别指明个人(’PER’)、组织(’ORG’)或地点(’LOC’)的特定命名实体。答案应遵循格式’实体名称, 实体类型’ 。
@@
When analyzing sentences in the China Securities Regulatory Commission filing documents, identify specific named entities that indicate individuals (’PER’), organizations (’ORG’), or locations (’LOC’). The answer should follow the format ’entity name, entity type’.
Relationship Extraction (FinRE)
请仔细分析所给金融报道和实体对,然后在[’合并’, ’竞争’,…]中选择能准确描述该实体对关系的选项。请直接给出答案,如有疑虑可回答unknown。
@@
Please carefully analyze the given financial report and entity pair, then choose the option from [’merge’, ’compete’, …] that accurately describes the relationship of the entity pair. Please provide the answer directly; if in doubt, you may respond with ’unknown’
Text Summarization (FinTS)
通过阅读金融公告,你的任务是对所给的文本进行简短的总结,重点突出主要论点,长度保持在一到两句之间。
@@
Your task is to provide a brief summary of the given text by reading financial announcements, emphasizing the main arguments, and keeping the length between one to two sentences
@@ F Evaluation Results of LLMs, FinLLMs and ICE-INTERN variants
Table 10: The detailed evaluation results of different 6B-7B LLMs, baseline FinLLMs and variants of ICE-INTERN on ICE-FLARE. All results via our evaluations are the average of three runs. Results with "*" of ChatGPT-3.5(ChatGPT) and ChatGPT-4(GPT4) on English dataset are cited from  (Xie et al., 2023b). Results in bold indicate the best results across all models. Results with "__" indicate the best results among models excluding ChatGPT.
@@ G Examples of Query Response among Various LLMs
Baichuan-7B ChatGPT-4 ICE-INTERN-full-7B
FinQA
浦东建设公告,公司子公司上海市浦东新区建设(集团)有限公司、上海浦兴路桥建设工程有限公司中标多项重大工程项目,中标金额总计为10.94亿元。请问中标事件对应的主体有哪些?
浦东建设、浦兴路桥、浦东建设(集团)
中标事件对应的主体有上海市浦东新区建设(集团)有限公司
上海市浦东新区建设(集团)有限公司,上海浦兴路桥建设工程有限公司
FinTS
市场普遍预期日本央行将维持其超宽松货币政策不变,但知情人士表示,周二将公布的经济展望报告中,可能上调4月起财政年度的通胀预估。能源和原材料成本的上涨已将批发价格推高至40年来的最高水平,企业已开始将上涨的成本转嫁给消费者。近期通胀温和上升,主要是投入成本上升和日元疲软推动的,而不是强劲的国内需求。日本央行行长黑田东彦曾表示,物价上涨应该伴随着工资的强劲增长。请对上述金融报道进行总结,给出具有中心主旨的摘要,回答要求尽量简短。
日本央行行长黑田东彦曾表示,物价上涨应该伴随着工资的强劲增长。
日本央行预计将上调4月起的通胀预估,尽管市场普遍预期货币政策将保持超宽松。成本上涨推高批发价格,企业已将涨成本转嫁给消费者,但通胀主要由投入成本上升和日元疲软推动,而非国内需求强劲。
知情人士:日本央行可能上调4月起财政年度的通胀预估
FinRE
新希望集团销售额约为上市公司<N>倍有余,由于背靠农业帝国“希望系”, 资本市场一直预期集团能够向新希望股份注入农业资产, 如果集团资产整合启动, 新希望股份的股。请分析上述涉及实体对(新希望股份,新希望集团)的关系。
被收购
被注资
自己
FinED
大智慧(601519)信披违规遭调查“互联网+”炒作或终结金鹰等多家基金公司高管离职 基民担忧业绩波动。请分析上述金融报道中的事件类型及及其事件主体,并以‘事件类型,事件主体’的格式输出。
套现约1.5亿元
交易违规, 大智慧
信批违规, 大智慧
FinER
本公司非执行董事孙月英女士不再担任中国远洋海运集团有限公司总会计师和中远财务有限责任公司董事长。请给出上述金融报道中存在的个人(’PER’)、组织(’ORG’)或地点(’LOC’)的特定命名实体,回答应遵循的格式’实体名称, 实体类型’。
中国远洋海运集团有限公司,ORG
中远海运发展股份有限公司, ORG
孙月英, PER; 中国远洋海运集团有限公司, ORG; 中远财务有限责任公司, ORG
FinNC
WTI原油涨幅回升至0.5%,现报75.58美元/桶。请对这个金融新闻报道进行分类,具体属于[’中国’,’外国’,’国际’,’公司’,’行业’,’大盘’,’经济’,’政策’,’政治’,’期货’,’债券’,’房地产’,’外汇’,’虚拟货币’,’新冠’,’能源’]中的哪些类别?。
输出与内容无关
国际,能源
国际期货

### v2

@@ E Examples of Instruction Prompt in Chinese Task
Sentiment Analysis (FinSA) 确定所提供的金融新闻文章中的句子的情绪,识别其中的情绪是积极,消极,还是中性的。只需回答’积极’、’中性’或’消极’。 Determine the sentiment of the sentences in the provided financial news article, identifying whether the sentiment is positive, negative, or neutral. Simply answer ’positive’, ’neutral’, or ’negative’.
@@
Sentiment Matching (FinSM) 判断给出的两个金融文本表达的意思是否相似,你只需要回答是或否。现在判断下面两个句子是否相似: Determine whether the two provided financial texts convey similar meanings; you only need to answer ’yes’ or ’no’. Now determine if the following two sentences are similar.
@@
News Classification (FinNC) 根据金融报道的内容,判断其属于’中国’、’国际’还是’外国’,只需给出其中一个类别。现在分析下面的报道,回答它所属的类别: According to the content of the financial report, determine whether it belongs to ’China’, ’International’, or ’Foreign’, only one category needs to be given. Now analyze the following report and answer the category it belongs to:
@@
Negative Judgment (FinNJ) 根据所给的金融新闻及实体,你需要判断所给实体是否含有负面的消息,你只需简单回答’有’或者’无’ Based on the given financial news and entities, you need to determine whether the given entities contain negative messages. You only need to simply answer ’yes’ or ’no’.
@@
Answer Selection (FinAS) 根据所提出的金融问题,从以下四个选项中选择最合适的一个。你的输出应该是:’A’、’B’、’C’ 或’D’。 Based on the financial issue presented, choose the most appropriate one from the following four options. Your output should be: ’A’, ’B’, ’C’, or ’D’.
@@
Stock Prediction (FinSP) 请仔细分析数据和推文,预测2017-10-13时\$trv的收盘价格是上涨还是下跌。请确认是上涨还是下跌, 只回答为’上涨’或者’下跌’。 Please carefully analyze the data and tweets, and predict whether the closing price of \$TRV on 2017-10-13 will go up or down. Please confirm whether it is going up or down. You can answer ’Rise’ or ’Fall’.
@@
Headline Classification (FinHC) 请考虑标题是否讨论了与黄金相关的过去事件。新闻标题是否暗示黄金的过去新闻?您的回答应该为’是’或’否’。 Please consider whether the headline discusses past events related to gold. Does the news headline imply past news about gold? Your answer should be ’Yes’ or ’No’.
@@
Question Answering (FinQA) 你需要分析金融文本,根据内容回答相关问题。如果你对答案感到迷茫,可以回答’无相应参数’。 You need to analyze financial texts and answer related questions based on the content. If you feel confused about the answer, you can reply with ’No relevant parameters’.
@@
Event Detection (FinED) 阅读以下金融领域的公告,判断所有的事件类型及其对应主体,省略数字。请以:’事件类型, 事件主体’的格式回复 。其中事件类型应该在这里面:[’信批违规’,…,’涉嫌欺诈’] The task is to read a financial announcement, determine all event types and their corresponding entities, and reply in the format ’event type, event subject’. The event types should be among those listed in the brackets [‘credit approval violation’,…,’suspected fraud’].
@@
Entity Recognition (FinER) 在分析中国证券交易委员会备案文件中的句子时,识别指明个人(’PER’)、组织(’ORG’)或地点(’LOC’)的特定命名实体。答案应遵循格式’实体名称, 实体类型’ 。 When analyzing sentences in the China Securities Regulatory Commission filing documents, identify specific named entities that indicate individuals (’PER’), organizations (’ORG’), or locations (’LOC’). The answer should follow the format ’entity name, entity type’.
@@
Relationship Extraction (FinRE) 请仔细分析所给金融报道和实体对,然后在[’合并’, ’竞争’,…]中选择能准确描述该实体对关系的选项。请直接给出答案,如有疑虑可回答unknown。 Please carefully analyze the given financial report and entity pair, then choose the option from [’merge’, ’compete’, …] that accurately describes the relationship of the entity pair. Please provide the answer directly; if in doubt, you may respond with ’unknown’
@@
Text Summarization (FinTS) 通过阅读金融公告,你的任务是对所给的文本进行简短的总结,重点突出主要论点,长度保持在一到两句之间。 Your task is to provide a brief summary of the given text by reading financial announcements, emphasizing the main arguments, and keeping the length between one to two sentences
@@ F Evaluation Results of LLMs, FinLLMs and ICE-INTERN variants
Table 11: The detailed evaluation results of different 6B-7B LLMs, baseline FinLLMs and variants of ICE-INTERN on ICE-FLARE. All results via our evaluations are the average of three runs. Results with "*" of ChatGPT(3.5-Trubo) and GPT-4 on English dataset are cited from  (Xie et al., 2023b). Results in bold indicate the best results across all models. Results with "__" indicate the best results among models excluding ChatGPT.
@@ G Examples of Query Response among Various LLMs
Baichuan-7B GPT-4 ICE-INTERN-full-7B
FinQA 浦东建设公告,公司子公司上海市浦东新区建设(集团)有限公司、上海浦兴路桥建设工程有限公司中标多项重大工程项目,中标金额总计为10.94亿元。请问中标事件对应的主体有哪些? 浦东建设、浦兴路桥、浦东建设(集团) 中标事件对应的主体有上海市浦东新区建设(集团)有限公司 上海市浦东新区建设(集团)有限公司,上海浦兴路桥建设工程有限公司
Pudong Construction announced that the company’s subsidiaries, Shanghai Pudong New Area Construction (Group) Company Limited and Shanghai Puxing Road and Bridge Construction Engineering Company Limited, have won a number of major projects, with a total winning amount of 1.094 billion yuan. What are the subjects corresponding to the winning bidding events? Pudong Construction, Puxing Road & Bridge, Pudong Construction (Group) The winning event corresponds to the subject of Shanghai Pudong New Area Construction (Group) Co. Shanghai Pudong New Area Construction (Group) Company Limited,Shanghai Puxing Road & Bridge Construction Engineering Co.
FinTS 市场普遍预期日本央行将维持其超宽松货币政策不变,但知情人士表示,周二将公布的经济展望报告中,可能上调4月起财政年度的通胀预估。能源和原材料成本的上涨已将批发价格推高至40年来的最高水平,企业已开始将上涨的成本转嫁给消费者。近期通胀温和上升,主要是投入成本上升和日元疲软推动的,而不是强劲的国内需求。日本央行行长黑田东彦曾表示,物价上涨应该伴随着工资的强劲增长。请对上述金融报道进行总结,给出具有中心主旨的摘要,回答要求尽量简短。 日本央行行长黑田东彦曾表示,物价上涨应该伴随着工资的强劲增长。 日本央行预计将上调4月起的通胀预估,尽管市场普遍预期货币政策将保持超宽松。成本上涨推高批发价格,企业已将涨成本转嫁给消费者,但通胀主要由投入成本上升和日元疲软推动,而非国内需求强劲。 知情人士:日本央行可能上调4月起财政年度的通胀预估。
The Bank of Japan is widely expected to leave its ultra-loose monetary policy unchanged, but people familiar with the matter said it may raise its inflation forecast for the fiscal year starting in April in its economic outlook report to be released on Tuesday. Rising energy and raw material costs have pushed wholesale prices to the highest level in 40 years, and companies have begun to pass on rising costs to consumers. The recent moderate rise in inflation has been driven mainly by higher input costs and a weaker yen, rather than strong domestic demand. Bank of Japan Governor Haruhiko Kuroda has said price increases should be accompanied by strong wage growth. Please summarize the above financial reports, giving a summary with the central thrust, and keep your answers as brief as possible. Bank of Japan Governor Haruhiko Kuroda has said that price increases should be accompanied by strong wage growth. The Bank of Japan is expected to raise its inflation forecast from April, although the market generally expects monetary policy to remain ultra-loose. Rising costs have pushed up wholesale prices and companies have passed on the cost increases to consumers, but inflation has been driven mainly by higher input costs and a weaker yen rather than strong domestic demand. BOJ may raise inflation forecast for fiscal year from April: people familiar with the matter.
FinRE 新希望集团销售额约为上市公司<N>倍有余,由于背靠农业帝国“希望系”, 资本市场一直预期集团能够向新希望股份注入农业资产, 如果集团资产整合启动, 新希望股份的股。请分析上述涉及实体对(新希望股份,新希望集团)的关系。 被收购 被注资 自己
New Hope Group’s sales are more than twice that of the listed company<N>, due to the backing of the agricultural empire "Hope", the capital market has been expecting the group to inject agricultural assets into New Hope shares, if the group’s asset integration is launched, the shares of New Hope shares. Please analyze the relationship between the entities involved in the above (New Hope, New Hope Group). Acquired Funded Self
FinED 大智慧(601519)信披违规遭调查“互联网+”炒作或终结金鹰等多家基金公司高管离职 基民担忧业绩波动。请分析上述金融报道中的事件类型及及其事件主体,并以‘事件类型,事件主体’的格式输出。 套现约1.5亿元 交易违规, 大智慧 信批违规, 大智慧
Great wisdom (601519) letter disclosure violations investigated "Internet +" speculation or the end of the Golden Eagle and many other fund company executives leave the funders worried about performance fluctuations. Please analyze the types of events and their event subjects in the above financial reports and output them in the format of ’event type, event subject’. Cashing in about \$150 million Trading Violations, Great Intelligence Letter Approval Violation, Great Intelligence
FinER 本公司非执行董事孙月英女士不再担任中国远洋海运集团有限公司总会计师和中远财务有限责任公司董事长。请给出上述金融报道中存在的个人(’PER’)、组织(’ORG’)或地点(’LOC’)的特定命名实体,回答应遵循的格式’实体名称, 实体类型’。 中国远洋海运集团有限公司,ORG 中远海运发展股份有限公司, ORG 孙月英, PER; 中国远洋海运集团有限公司, ORG; 中远财务有限责任公司, ORG
Ms. Sun Yueying, a non-executive director of the Company, ceased to be the chief accountant of China Ocean Shipping Group Company Limited and the chairman of COSCO Finance Company Limited. Please give the name of the person (’PER’), organization (’ORG’) or location (’LOC’) of the particular named entity that exists in the above financial report. ) of a specific named entity, the response should follow the format ’Entity Name, Entity Type’. China Ocean Shipping Group Company Limited, ORG COSCO Shipping Development Company Limited, ORG Sun Yueying, PER; China Ocean Shipping Group Company Limited, ORG; COSCO Finance LLC, ORG.
FinNC WTI原油涨幅回升至0.5%,现报75.58美元/桶。请对这个金融新闻报道进行分类,具体属于[’中国’,’外国’,’国际’,’公司’,’行业’,’大盘’,’经济’,’政策’,’政治’,’期货’,’债券’,’房地产’,’外汇’,’虚拟货币’,’新冠’,’能源’]中的哪些类别?。 输出与内容无关 国际,能源 国际期货
WTI crude oil has rallied to 0.5% and is now at \$75.58 per barrel. Please categorize this financial news report, specifically which categories in [’China’,’Foreign’,’International’,’Companies’,’Industries’,’Broader Markets’,’Economy’,’Policies’,’Politics’,’Futures’,’Bonds’,’Real Estate’,’Forex’,’Virtual Currencies’,’New Crowns’,’Energy’]. output independent of content International, Energy International futures